OpenStudy (anonymous):

1/9(y+18)+1/3(2y+3)=y+3 step by step plzz

OpenStudy (anonymous):

multiply both sides by 9 to clear the annoying fractions

OpenStudy (anonymous):

i.e. start with \[y+18+3(2y+3)=9(y+3)\] as a first step

OpenStudy (anonymous):

that was wrong \[y+18+6y+9=9y+27\] \[7y+27=9y+27\] \[7y=9y\] \[y=0\]